OpenStudy (anonymous):

Solve the system equations . 3x - 4y = -10 and 5x + 8y = -2 . A. (1,2) B. (2,1) C. (1,-2) D. (-2,1) ?

myininaya (myininaya):

3x-4y=-10 5x+8y=-2 Multiply first equation by 2 6x-8y=-20 5x+8y=-2 Now add these equations! 11x+0=-22 11x=-22 .... Will you show me how to solve the rest of this?

OpenStudy (anonymous):

Just a moment

OpenStudy (anonymous):

x = -2

myininaya (myininaya):

Yep Yep! :)

myininaya (myininaya):

\[3(-2)-4y=-10\] I just replaced x with -2 in one of the equations do you think you can solve for y

OpenStudy (anonymous):

y = -4 ?

myininaya (myininaya):

-6-4y=-10 +6 +6 -4y=-4 y=1

myininaya (myininaya):

nice try though alexis! :)